OpenStudy (anonymous):

find the equation of the line containing (3,1) and (9,5)

OpenStudy (anonymous):

Find the slope by taking the rise over the run. The rise is 5-1=4 The run is 9-3=6 So the slope is 4/6 Enter the slope in the equation of a line. So you have this y=4/6x+b Substitute one of the points into the equation. If I pick the first point (3,1) I have this 1=4/6(3)+b Solve for b So b= 1 So the equation is y=2/3 + 1

OpenStudy (anonymous):

Sorry b should -1

OpenStudy (anonymous):

Here's a graph of the equation y=2/3x-1